OpenStudy (anonymous):

What is the value of the expression 2/4^-3 ? A. 1/32 B. 1/6 C. 128 D. 256

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

@imqwerty

imqwerty (imqwerty):

\[4^\left( -3 \right)=\frac{ 1 }{ 4^3 } \] \[\frac{ 1 }{ 4^\left( -3 \right) }=4^3\] multiply 2 on both sides \[\frac{ 2 }{ 4^\left( -3 \right) }=2 \times 4^3=128\]